# ViEEG: Hierarchical Neural Coding with Cross-Modal Progressive Enhancement for EEG-Based Visual Decoding

### v2

@@ 4.3. Hierarchical EEG encoding
Our neuro-inspired encoding architecture extracts hierarchical (Contour/object/context) EEG features. Given raw EEG signals $E$, three parallel streams extract cortical hierarchy-aligned EEG features through spatiotemporal convolution (STConv) and cross-attention hierarchical integration (CAHI).
4.4. Spatiotemporal Convolution
@@
[…]
4.5. Cross-Attention Hierarchical Integration
The CAHI integrates biologically inspired features through contour-to-object and object-to-context feature integration. Since contour features can enhance existing object features, and object features provide additional context information, we integrate these two integration processes accordingly. Assuming that the low-level representation is denoted by $a$ and the higher-level representation (object) by $b$, we perform bottom-up information integration, i.e., we transfer information from $a$ to $b$. We use $a$ and apply several linear transformations to generate the key embeddings for the attention mechanism, while $b$ is used to generate the query and value embeddings:
(6) $K^{i}=a^{i}W^{K},\quad Q^{i}=b^{i}W^{Q},\quad V^{i}=a^{i}W^{V}$
@@ 4.5. Cross-Attention Hierarchical Integration
(7) $F^{i}=\text{Attention}(Q^{i},K^{i},V^{i})=\text{softmax}\left(\frac{Q^{i}{K^{i}}^{T}}{\sqrt{d}}\right)V^{i}$
@@
(8) $F=\text{MHA}(K=a,Q=V=b)=(F^{1}\|F^{2}\|\dots\|F^{h})W^{O}+b$
[…]
Contour-to-object Integration: The low-level representation is the contour feature $F_{b}^{init}$, while the high-level representation is the object feature $F_{f}^{init}$. The refined object feature is computed as:
(9) $F_{f}=\text{LN}\Big{(}F_{f}^{init}+\text{Dropout}\Big{(}\text{MHA}(K=F_{b}^{init},Q=V=F_{f}^{init})\Big{)}\Big{)}$
Object-to-context Integration: The low-level representation is the object feature $F_{f}^{init}$, while the high-level representation is the context feature $F_{r}^{init}$. The refined context feature is computed as:
(10) $F_{r}=\text{LN}\Big{(}F_{r}^{init}+\text{Dropout}\Big{(}\text{MHA}(K=F_{f},Q=V=F_{r}^{init})\Big{)}\Big{)}$
4.6. Hierarchical Contrastive Learning
[…]
(11) $\mathcal{L}=\text{CrossEntropy}\left(\alpha\cdot\cos(F_{\text{eeg}},C_{\text{img}}),\,Y\right)$
@@ 5. Experiment
BraVL (Du et al., 2023) 6.1 17.9 4.9 14.9 5.6 17.4 5.0 15.1 4.0 13.4 6.0 18.2 6.5 20.4 8.8 23.7 4.3 14.0 7.0 19.7 5.8 17.5
LSTM (Hochreiter, 1997) 16.4 43.5 16.1 40.2 20.7 49.4 18.6 42.3 14.2 35.2 18.2 43.1 18.7 47.2 22.0 47.9 16.8 43.8 21.7 52.2 18.3 44.5
ConvNet (Schirrmeister et al., 2017) 14.9 39.3 18.9 42.1 17.8 49.0 23.9 55.8 12.2 32.9 20.1 46.5 15.5 42.6 20.7 48.9 20.7 49.8 19.3 47.5 18.4 45.5
Conformer (Song et al., 2022) 11.4 32.4 15.2 41.9 19.8 50.9 23.0 56.6 13.6 33.4 18.1 49.0 18.5 48.2 27.1 56.9 15.2 40.0 22.6 57.8 18.5 46.7
DGCNN (Song et al., 2018) 12.3 36.6 11.5 39.5 15.7 43.8 19.7 50.6 10.6 32.6 15.4 46.6 14.0 43.2 25.1 54.5 16.0 43.7 17.9 53.0 15.8 44.4
EEGNet (Lawhern et al., 2018) 16.0 42.9 17.9 48.6 18.2 51.5 23.9 59.0 14.4 37.7 19.5 52.0 18.5 50.2 30.2 61.2 23.3 51.2 22.5 58.3 20.4 51.2
Mb2C (Wei et al., 2024) 23.6 56.3 22.6 50.5 26.3 60.1 34.8 67.0 21.3 53.0 31.0 62.3 25.0 54.8 39.0 69.3 27.5 59.3 33.1 70.8 28.4 60.3
NICE (Song et al., 2024) 21.7 51.2 23.3 55.0 29.1 60.5 32.3 69.6 18.2 45.6 29.3 62.1 24.3 59.2 41.3 72.4 24.3 59.0 28.9 62.6 27.3 59.7
ATM (Li et al., 2024) 25.6 60.5 22.0 54.5 25.0 62.4 31.4 60.9 12.9 43.0 21.3 51.1 30.5 61.5 38.8 72.0 24.4 51.5 29.1 63.5 26.1 58.1
CogCap (Zhang et al., 2024) 27.2 59.5 28.7 56.9 37.1 66.1 37.6 63.2 21.8 47.7 31.5 58.0 32.8 59.5 47.6 73.5 33.3 57.6 35.0 63.5 33.3 60.5
@@
BraVL (Du et al., 2023) 2.3 8.0 1.5 6.3 1.4 5.9 1.7 6.7 1.5 5.6 1.8 7.2 2.1 8.1 2.2 7.6 1.6 6.4 2.3 8.5 1.8 7.0
LSTM (Hochreiter, 1997) 9.2 30.3 10.3 26.4 6.4 19.3 9.1 25.2 7.8 22.2 8.4 27.6 7.5 22.0 10.6 26.9 5.3 22.6 9.3 26.1 8.4 24.8
ConvNet (Schirrmeister et al., 2017) 10.4 31.5 14.2 35.0 8.7 26.9 12.1 31.4 5.7 22.5 10.2 28.0 8.3 22.5 10.0 29.9 6.6 19.8 12.2 34.0 9.8 28.2
Conformer (Song et al., 2022) 6.3 22.3 5.7 20.4 5.8 15.7 7.8 21.8 6.7 18.4 10.4 32.9 7.0 24.1 9.1 25.2 5.0 17.2 11.7 33.2 7.5 23.1
DGCNN (Song et al., 2018) 10.8 30.6 11.9 31.3 6.3 21.3 8.2 24.6 6.1 18.5 11.2 30.4 6.8 20.5 11.0 28.6 9.4 25.4 12.2 30.6 9.4 26.2
EEGNet (Lawhern et al., 2018) 11.0 31.1 11.4 34.0 6.5 24.7 13.4 34.4 6.7 26.7 8.5 29.4 7.9 22.0 10.7 32.9 8.9 27.7 15.1 41.6 10.0 30.4
Mb2C (Wei et al., 2024) 10.5 28.1 11.3 32.8 8.8 27.6 13.6 33.5 10.6 27.5 12.1 33.1 11.5 31.8 12.0 32.1 12.1 31.3 16.1 42.1 11.9 32.0
NICE (Song et al., 2024) 13.5 39.3 16.7 42.7 12.4 35.5 17.9 41.9 14.7 38.1 15.6 44.6 13.6 39.0 13.5 37.2 17.0 42.0 22.8 50.7 15.7 41.1
ATM (Li et al., 2024) 17.1 41.8 20.2 44.2 13.2 36.7 17.0 40.7 15.1 41.0 13.5 38.3 10.1 29.0 15.2 41.9 13.5 38.4 20.0 45.4 15.5 39.6
CogCap (Zhang et al., 2024) 16.3 42.3 16.2 37.9 8.8 26.8 15.4 37.6 10.1 31.7 14.0 35.4 10.7 26.9 13.9 34.2 9.0 32.4 15.3 38.6 13.0 34.4
@@
Table 2. Overall accuracy (%) comparison: Top-1 and Top-5 in 200-way zero-shot object recognition
@@
$F_{b}$-Only 14.2 41.8 14.0 39.8 17.0 39.1 20.9 53.2 8.6 30.3 13.6 40.9 14.7 40.7 17.6 46.4 12.8 37.6 14.8 44.0 14.8 41.4
$F_{f}$-Only 22.2 52.1 24.5 56.1 29.9 66.0 34.0 67.1 19.2 50.0 32.5 63.5 26.5 58.7 38.7 73.9 29.3 61.8 31.8 68.5 28.9 61.8
$F_{r}$-Only 22.1 52.6 24.8 58.0 32.4 62.7 33.1 67.4 18.5 49.1 29.7 66.2 26.8 60.1 40.6 74.1 27.4 62.5 31.8 68.6 28.6 62.1
w/o C-Att 32.4 66.8 34.6 65.9 39.1 71.3 46.5 80.2 27.2 59.8 40.5 74.3 35.3 71.3 52.2 80.8 34.6 71.9 43.9 75.5 38.6 71.8
@@
$F_{b}$-Only 5.9 18.6 7.9 22.7 5.7 19.1 7.4 26.3 7.6 20.5 4.2 18.6 9.0 23.1 7.9 21.8 7.0 20.8 10.2 30.9 7.3 22.3
$F_{f}$-Only 15.8 43.0 19.8 42.5 13.3 40.3 20.5 45.1 13.7 37.1 17.1 41.7 13.2 39.7 15.5 41.9 15.3 42.9 23.4 51.4 16.8 42.6
$F_{r}$-Only 14.9 41.5 18.4 45.9 14.1 41.3 18.1 42.2 15.9 38.3 18.5 42.8 15.4 41.5 16.0 41.6 18.3 45.3 20.5 50.2 17.0 43.1
w/o C-Att 19.1 52.1 22.0 48.9 15.1 45.1 24.2 55.2 16.7 41.6 21.0 50.4 17.2 44.7 19.5 45.6 20.9 48.8 31.1 59.8 20.7 49.2
@@
Table 3. Ablation study: Top-1 and Top-5 accuracy (%) in 200-way zero-shot object recognition
5.1. Dataset
The THINGS-EEG dataset (Gifford et al., 2022) is a large-scale benchmark for zero-shot EEG-based object recognition, comprising brain responses from ten participants exposed to images from the THINGS database under a rapid serial visual presentation (RSVP) (Thorpe et al., 1996) paradigm. Each stimulus was displayed for 100 ms, followed by a 100 ms inter-stimulus interval (ISI) with a blank screen. The dataset comprises 1654 training concepts and 200 zero-shot test concepts, yielding a total of 82,160 trials per subject. EEG signals were acquired using a 64-channel cap, band-pass filtered between 0.1–100 Hz, and subsequently downsampled to 100 Hz. Preprocessing involved segmenting epochs from 200 ms before to 800 ms after stimulus onset, followed by baseline correction using the pre-stimulus interval.
5.2. Experimental Details
@@
For fair comparison, experiments were conducted under subject-dependent and subject-independent settings, following the protocol of ATM (Li et al., 2024). All methods used the same CLIP backbone (CLIP-ViT-H/14777https://huggingface.co/laion/CLIP-ViT-bigG-14-laion2B-39B-b160k) to ensure alignment consistency. For subject-dependent training, 740 trials were randomly selected as the validation set, and the model with the lowest validation loss was retained. In the subject-independent setting, we employed a leave-one-subject-out (LOSO) protocol, using 6660 trials for validation in each fold. All experiments were repeated five times, and the average test accuracy was reported to mitigate variance and ensure result robustness.
Metric ViEEG NICE Improve Best Case (Sub.10)
Avg. Top-1 (Dep.) 40.9% 27.3% 49.82% 42.8% vs 28.9% ($\uparrow$ 48.1%)
Avg. Top-5 (Dep.) 74.5% 59.7% 24.79% 79.8% vs 62.6% ($\uparrow$ 27.5%)
Avg. Top-1 (Ind.) 22.9% 15.7% 45.86% 31.2% vs 22.8% ($\uparrow$ 36.8%)
Avg. Top-5 (Ind.) 51.4% 41.1% 25.06% 60.3% vs 50.7% ($\uparrow$ 19.0%)
Table 4. Summary of recognition performance and relative improvement over NICE across Top-1 and Top-5 metrics.
5.3. Overall Performance
@@
The experimental comparison against state-of-the-art methods demonstrates ViEEG’s superior decoding capability across both evaluation paradigms. We conduct comprehensive comparisons against 10 EEG decoding methods, including BraVL (Du et al., 2023), LSTM (Hochreiter, 1997), ConvNet (Schirrmeister et al., 2017), Conformer (Song et al., 2022), DGCNN (Song et al., 2018), EEGNet (Lawhern et al., 2018), MB2C (Wei et al., 2024), NICE (Song et al., 2024), ATM (Li et al., 2024), and CognitionCapturer (CogCap) (Zhang et al., 2024). Table 2 summarizes the 200-way zero-shot recognition performance across both experimental paradigms, and our method ViEEG achieves consistent improvements over existing approaches through biologically grounded hierarchical modeling.
@@
Subject-dependent Experiments: ViEEG achieves new performance benchmarks, reaching 40.9% Top-1 and 74.5% Top-5 average accuracy across 10 subjects, outperforming the previous SOTA method by substantial margins. Specifically, compared to the baseline NICE, ViEEG demonstrates an average improvement of 49.82% in Top-1 accuracy and 24.79% in Top-5 accuracy, reflecting its capacity to effectively exploit subject-specific neural patterns. This performance highlights its enhanced capacity of ViEEG to handle visual scenes through hierarchical feature disentanglement.
@@
Cross-subject Experiments: When tested on unseen subjects under LOSO protocol, ViEEG maintains robust performance with 22.9% Top-1 and 51.4% Top-5 accuracy. The biological plausibility of our hierarchical design enables better generalization compared to NICE’s flat representation approach. For average object recognition accuracy, ViEEG demonstrates relative improvement of 45.86% in Top-1 accuracy and 25.06% in Top-5 accuracy. This performance gap stems from ViEEG’s hierarchical feature disentanglement that captures invariant neural patterns across individuals.
Consistent performance gains across all evaluation metrics (Table 4) statistically validate that explicit modeling of visual processing hierarchy substantially improves EEG decoding accuracy.
5.4. ablation Study

### v3

@@ 3.4 Hierarchical EEG Encoding
Our neuro-inspired encoding architecture extracts hierarchical (contour/object/context) EEG features. Given raw EEG $E$, three parallel streams extract cortical hierarchy-aligned EEG features through spatiotemporal convolution (STConv) and cross-attention hierarchical integration (CAHI).
3.5 Spatiotemporal Convolution
@@
$F^{(1)}$ $=\text{ELU}(\text{BatchNorm}(\text{AvgPool}(\text{Conv2D}_{t}(E))))$ (5)
$F^{(2)}$ $=\text{ELU}(\text{BatchNorm}(\text{Conv2D}_{s}(F^{(1)})))$
$F^{0}$ $=\text{Conv2D}_{proj}(F^{(2)})$
where $\text{Conv2D}_{t}(\cdot)$ is temporal convolution, $\text{Conv2D}_{s}(\cdot)$ represents spatial convolution, and $\text{Conv2D}_{proj}(\cdot)$ is a $1\times 1$ convolution. These hierarchical spatiotemporal features facilitate robust representation learning, which is further refined through the cross-attention mechanism. After three parallel STConv operations, EEG signals are decoupled into three features: $F_{b}^{0}$, $F_{f}^{0}$, and $F_{r}^{0}$.
3.6 Cross-Attention Hierarchical Integration (CAHI)
The CAHI integrates biologically inspired features through contour-to-object and object-to-context integration. Since contour features can enhance object features, and object features provide additional context information, we integrate these two integration processes accordingly. Assuming that low-level representation is denoted by $a$ and higher-level representation by $b$, we perform bottom-up integration, and transfer information from $a$ to $b$. We use $a$ and apply several linear transformations to generate key embeddings in attention, while generate query and value embeddings from $b$ as:
$K^{i}=a^{i}W^{K},\quad Q^{i}=b^{i}W^{Q},\quad V^{i}=a^{i}W^{V}$ (6)
@@ 3.6 Cross-Attention Hierarchical Integration (CAHI)
$F^{i}=\text{Attention}(Q^{i},K^{i},V^{i})=\text{softmax}\left(\frac{Q^{i}{K^{i}}^{T}}{\sqrt{d}}\right)V^{i}$ (7)
@@
$F=\text{MHA}(K=a,Q=V=b)=(F^{1}\|F^{2}\|\dots\|F^{h})W^{O}+b$ (8)
where $\text{MHA}(\cdot)$ denote multi-head attention function, $h$ is the number of attention heads, $W^{O}$ is a linear transformation, and $\|$ denotes concatenation.
Contour-to-Object Integration: The low-level and high-level representation is contour feature $F_{b}^{0}$ and object feature $F_{f}^{0}$, respectively, and refined object feature is computed as:
$F_{f}=\text{LN}\Big{(}F_{f}^{0}+\text{Dropout}\Big{(}\text{MHA}(K=F_{b}^{0},Q=V=F_{f}^{0})\Big{)}\Big{)}$ (9)
where $\text{LN}(\cdot)$ is the LayerNorm function.
Object-to-Context Integration: The low-level and high-level representation is object feature $F_{f}^{0}$ and context feature $F_{r}^{0}$, respectively, and refined context feature is computed as:
$F_{r}=\text{LN}\Big{(}F_{r}^{0}+\text{Dropout}\Big{(}\text{MHA}(K=F_{f},Q=V=F_{r}^{0})\Big{)}\Big{)}$ (10)
3.7 Hierarchical Contrastive Learning
To further enhance representation learning, we adopt a contrastive learning framework that aligns the concatenated EEG features with their corresponding image embeddings. Specifically, we first concatenate the hierarchical EEG features, i.e., $F_{\text{eeg}}=(F_{b}\|F_{f}\|F_{r})$, and similarly, the image features are concatenated as $C_{\text{img}}=(C_{b}\|C_{f}\|C_{r})$. We then compute the cosine similarity between these concatenated features, scaled by a learnable temperature parameter $\alpha$ (with $\alpha=\exp(\text{logit\_scale})$). The final loss is computed as the cross-entropy loss on the resulting similarity logits:
$\mathcal{L}=\text{CrossEntropy}\left(\alpha\cdot\cos(F_{\text{eeg}},C_{\text{img}}),\,Y\right)$ (11)
@@ 4 Experiment
BraVL (Du et al. 2023) 6.1 17.9 4.9 14.9 5.6 17.4 5.0 15.1 4.0 13.4 6.0 18.2 6.5 20.4 8.8 23.7 4.3 14.0 7.0 19.7 5.8 17.5
LSTM (Hochreiter 1997) 16.4 43.5 16.1 40.2 20.7 49.4 18.6 42.3 14.2 35.2 18.2 43.1 18.7 47.2 22.0 47.9 16.8 43.8 21.7 52.2 18.3 44.5
ConvNet (Schirrmeister et al. 2017) 14.9 39.3 18.9 42.1 17.8 49.0 23.9 55.8 12.2 32.9 20.1 46.5 15.5 42.6 20.7 48.9 20.7 49.8 19.3 47.5 18.4 45.5
Conformer (Song et al. 2022) 11.4 32.4 15.2 41.9 19.8 50.9 23.0 56.6 13.6 33.4 18.1 49.0 18.5 48.2 27.1 56.9 15.2 40.0 22.6 57.8 18.5 46.7
DGCNN (Song et al. 2018) 12.3 36.6 11.5 39.5 15.7 43.8 19.7 50.6 10.6 32.6 15.4 46.6 14.0 43.2 25.1 54.5 16.0 43.7 17.9 53.0 15.8 44.4
EEGNet (Lawhern et al. 2018) 16.0 42.9 17.9 48.6 18.2 51.5 23.9 59.0 14.4 37.7 19.5 52.0 18.5 50.2 30.2 61.2 23.3 51.2 22.5 58.3 20.4 51.2
Mb2C (Wei et al. 2024) 23.6 56.3 22.6 50.5 26.3 60.1 34.8 67.0 21.3 53.0 31.0 62.3 25.0 54.8 39.0 69.3 27.5 59.3 33.1 70.8 28.4 60.3
NICE (Song et al. 2024) 21.7 51.2 23.3 55.0 29.1 60.5 32.3 69.6 18.2 45.6 29.3 62.1 24.3 59.2 41.3 72.4 24.3 59.0 28.9 62.6 27.3 59.7
ATM (Li et al. 2024) 25.6 60.5 22.0 54.5 25.0 62.4 31.4 60.9 12.9 43.0 21.3 51.1 30.5 61.5 38.8 72.0 24.4 51.5 29.1 63.5 26.1 58.1
CogCap (Zhang et al. 2024) 27.2 59.5 28.7 56.9 37.1 66.1 37.6 63.2 21.8 47.7 31.5 58.0 32.8 59.5 47.6 73.5 33.3 57.6 35.0 63.5 33.3 60.5
@@
BraVL (Du et al. 2023) 2.3 8.0 1.5 6.3 1.4 5.9 1.7 6.7 1.5 5.6 1.8 7.2 2.1 8.1 2.2 7.6 1.6 6.4 2.3 8.5 1.8 7.0
LSTM (Hochreiter 1997) 9.2 30.3 10.3 26.4 6.4 19.3 9.1 25.2 7.8 22.2 8.4 27.6 7.5 22.0 10.6 26.9 5.3 22.6 9.3 26.1 8.4 24.8
ConvNet (Schirrmeister et al. 2017) 10.4 31.5 14.2 35.0 8.7 26.9 12.1 31.4 5.7 22.5 10.2 28.0 8.3 22.5 10.0 29.9 6.6 19.8 12.2 34.0 9.8 28.2
Conformer (Song et al. 2022) 6.3 22.3 5.7 20.4 5.8 15.7 7.8 21.8 6.7 18.4 10.4 32.9 7.0 24.1 9.1 25.2 5.0 17.2 11.7 33.2 7.5 23.1
DGCNN (Song et al. 2018) 10.8 30.6 11.9 31.3 6.3 21.3 8.2 24.6 6.1 18.5 11.2 30.4 6.8 20.5 11.0 28.6 9.4 25.4 12.2 30.6 9.4 26.2
EEGNet (Lawhern et al. 2018) 11.0 31.1 11.4 34.0 6.5 24.7 13.4 34.4 6.7 26.7 8.5 29.4 7.9 22.0 10.7 32.9 8.9 27.7 15.1 41.6 10.0 30.4
Mb2C (Wei et al. 2024) 10.5 28.1 11.3 32.8 8.8 27.6 13.6 33.5 10.6 27.5 12.1 33.1 11.5 31.8 12.0 32.1 12.1 31.3 16.1 42.1 11.9 32.0
NICE (Song et al. 2024) 13.5 39.3 16.7 42.7 12.4 35.5 17.9 41.9 14.7 38.1 15.6 44.6 13.6 39.0 13.5 37.2 17.0 42.0 22.8 50.7 15.7 41.1
ATM (Li et al. 2024) 17.1 41.8 20.2 44.2 13.2 36.7 17.0 40.7 15.1 41.0 13.5 38.3 10.1 29.0 15.2 41.9 13.5 38.4 20.0 45.4 15.5 39.6
CogCap (Zhang et al. 2024) 16.3 42.3 16.2 37.9 8.8 26.8 15.4 37.6 10.1 31.7 14.0 35.4 10.7 26.9 13.9 34.2 9.0 32.4 15.3 38.6 13.0 34.4
@@
Table 2: Overall accuracy (%) comparison: Top-1 and Top-5 in 200-way zero-shot object recognition.
@@
$F_{b}$ Feature Only 14.2 41.8 14.0 39.8 17.0 39.1 20.9 53.2 8.6 30.3 13.6 40.9 14.7 40.7 17.6 46.4 12.8 37.6 14.8 44.0 14.8 41.4
$F_{f}$ Feature Only 22.2 52.1 24.5 56.1 29.9 66.0 34.0 67.1 19.2 50.0 32.5 63.5 26.5 58.7 38.7 73.9 29.3 61.8 31.8 68.5 28.9 61.8
$F_{r}$ Feature Only 22.1 52.6 24.8 58.0 32.4 62.7 33.1 67.4 18.5 49.1 29.7 66.2 26.8 60.1 40.6 74.1 27.4 62.5 31.8 68.6 28.6 62.1
w/o Cross-Attention 32.4 66.8 34.6 65.9 39.1 71.3 46.5 80.2 27.2 59.8 40.5 74.3 35.3 71.3 52.2 80.8 34.6 71.9 43.9 75.5 38.6 71.8
@@
$F_{b}$ Feature Only 5.9 18.6 7.9 22.7 5.7 19.1 7.4 26.3 7.6 20.5 4.2 18.6 9.0 23.1 7.9 21.8 7.0 20.8 10.2 30.9 7.3 22.3
$F_{f}$ Feature Only 15.8 43.0 19.8 42.5 13.3 40.3 20.5 45.1 13.7 37.1 17.1 41.7 13.2 39.7 15.5 41.9 15.3 42.9 23.4 51.4 16.8 42.6
$F_{r}$ Feature Only 14.9 41.5 18.4 45.9 14.1 41.3 18.1 42.2 15.9 38.3 18.5 42.8 15.4 41.5 16.0 41.6 18.3 45.3 20.5 50.2 17.0 43.1
w/o Cross-Attention 19.1 52.1 22.0 48.9 15.1 45.1 24.2 55.2 16.7 41.6 21.0 50.4 17.2 44.7 19.5 45.6 20.9 48.8 31.1 59.8 20.7 49.2
@@
Table 3: Ablation study: Top-1 and Top-5 accuracy (%) in 200-way zero-shot object recognition.
4.1 Dataset
The THINGS-EEG dataset (Gifford et al. 2022) is the large-scale benchmark and the most widely used in zero-shot EEG object recognition, comprising brain responses from ten participants exposed to images from the THINGS database under a RSVP paradigm. The dataset comprises 1654 training concepts and 200 zero-shot test concepts, yielding a total of 82,160 trials per subject. EEG signals were acquired using a 64-channel cap, band-pass filtered between 0.1–100 Hz, and subsequently downsampled to 100 Hz. Preprocessing involved segmenting epochs from 200 ms before to 800 ms after stimulus onset, followed by baseline correction using the pre-stimulus interval. Additionally, we evaluate ViEEG on THINGS-MEG dataset (Hebart et al. 2023) in Appendix, to validate the universality across various neural signals.
4.2 Experimental Details
@@
For fair comparison, experiments were conducted follow the protocol of NICE/ATM. All methods used the same CLIP backbone (CLIP-ViT-H/14) to ensure alignment consistency. For subject-dependent training, 740 trials were randomly selected as the validation set, and the model with the lowest validation loss was retained. For subject-independent training, we employed a leave-one-subject-out (LOSO) protocol, using 6660 trials for validation in each fold. All experiments were repeated five times, and the average test accuracy was reported to mitigate variance and ensure result robustness. Importantly, additional experimental details, including THINGS-MEG results, extensive ablation and parameter studies, computational resource analysis, visualizations, and more image retrieval and reconstruction examples, are provided in supplementary Appendix.
4.3 Overall Performance
@@
The experimental comparison against state-of-the-art methods demonstrates ViEEG’s superior decoding capability across both evaluation paradigms. We conduct comprehensive comparisons against 10 EEG decoding methods, including BraVL (Du et al. 2023), LSTM (Hochreiter 1997), ConvNet (Schirrmeister et al. 2017), Conformer (Song et al. 2022), DGCNN (Song et al. 2018), EEGNet (Lawhern et al. 2018), MB2C (Wei et al. 2024), NICE (Song et al. 2024), ATM (Li et al. 2024), and CognitionCapturer (CogCap) (Zhang et al. 2024). Table 2 summarizes the 200-way zero-shot recognition performance across both experimental paradigms, and our method ViEEG achieves consistent improvements over existing approaches through biologically grounded hierarchical modeling.
@@
Subject-Dependent Experiment. ViEEG achieves new performance benchmarks, reaching 40.9% Top-1 and 74.5% Top-5 average accuracy across 10 subjects, outperforming the previous SOTA method by substantial margins. Specifically, compared to the baseline NICE, ViEEG demonstrates an average improvement of 49.82% in Top-1 accuracy and 24.79% in Top-5 accuracy, reflecting its capacity to effectively exploit subject-specific neural patterns. This performance highlights ViEEG’s capacity to handle visual scenes through hierarchical feature disentanglement.
@@
Subject-Independent Experiment. When tested on unseen subjects under LOSO protocol, ViEEG maintains robust performance with 22.9% Top-1 and 51.4% Top-5 accuracy. The biological plausibility of hierarchical design enables better generalization compared to flat representation. For average object recognition accuracy, ViEEG demonstrates relative improvement of 45.86% in Top-1 accuracy and 25.06% in Top-5 accuracy. This performance gap stems from ViEEG’s hierarchical feature disentanglement that captures invariant neural patterns across individuals.
4.4 Ablation Study